Circular No. 6210

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-1 W. Paciesas, D. Crary, G. Fishman, B. A. Harmon, C. Kouveliotou,

and S. N. Zhang, Marshall Space Flight Center, NASA, report for the Compton Observatory BATSE team: “Since about June 28, the blackholecandidate binary system Cyg X-1 (= HDE 226868) has been in a state of unusually intense low-frequency noise. The power spectrum of the 20- to 100-keV emission between 0.005 and 0.5 Hz shows the typical shape for Cyg X-1: flat at low frequencies and falling off roughly as 1/f at high frequencies. However, the total power in the low-frequency noise increased by a factor of ∼ 3 between June 28 and July 30, while the break frequency dropped from ∼ 0.08 to ∼ 0.04 Hz, similar to the effect seen at lower energies by Belloni and Hasinger (1990, A.Ap. 227, L33). The increased noise is not correlated with source intensity. Individual Cyg X-1 flares with timescales of order 1–10 s have triggered the burst detection algorithm aboard BATSE > 30 times since June 28. Continuous BATSE monitoring has shown no similar level of activity in Cyg X-1 since 1991 Nov.”

X-RAY NOVA 1994 IN SCORPIUS H. Inoue, F. Nagase, and Y. Ueda report on behalf of the ASCA team:

“ASCA observed GRO J1655–40 during Aug. 15.45–16.18 UT. The average flux in the range 2–10 keV was 3.4 × 10^{−8} erg cm^{−2} s^{−1} (∼ 1.6 Crab, remaining within ± 10 percent of this throughout), and was ∼ 3 times as bright as that in the previous ASCA observation in 1994 Sept. (IAUC 6094). The preliminary spectral analysis shows that the 1- to 10-keV spectrum can be reproduced by a power law with an exponential cutoff and with a photoelectric absorption. The best-fit photon index, folding energy, and absorption column are 0.5, 2 keV, and 8 × 10^{21} cm^{−2}, respectively.”

COMET 41P/TUTTLE-GIACOBINI-KRESÁK Total visual magnitude estimates (cf. IAUC 6207): Aug. 19.36 UT, 7.7

(D. Seargent, The Entrance, N.S.W., 25×100 binoculars); 20.17, 8.3 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector; tails 8′ and 4′ long in p.a. 110^{o} and 335^{o}, and faint northward fan of connecting material in between); 20.93, 8.9 (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars).

1995 August 22 Daniel W. E. Green